# Sketch2PoseNet: Efficient and Generalized Sketch to 3D Human Pose Prediction

### v1

@@ 4.2. Sketch Feature Extractor
It is also known that ControlNet leverages trainable copies of the encoding layers within the denoising U-Net, serving as a robust backbone to learn various conditional controls, significantly enhancing the fine-grained spatial controllability of the Latent Diffusion Model (LDM)  (Rombach et al., 2022). In our implementation, we utilize the ControlNet architecture to handle pose-conditioned information from the 2D guidance $\mathcal{G}$ and integrate it into the image features within the denoising U-Net $\boldsymbol{\epsilon_{\theta}}$. The output $\mathcal{F}$ in the decoding layers of $\boldsymbol{\epsilon_{\theta}}$ is expressed as:
@@ 4.3. SMPL Mesh Regressor
In the last stage, an SMPL mesh regressor is proposed to predict 3D poses from the previously extracted features $\mathcal{F}$. Specifically, $p_{\phi_{3}}(\boldsymbol{y}\!\mid\!\mathcal{F})$ refers to the prediction head that generates parameters of the body model from the hierarchical feature maps $\mathcal{F}$. We first lift the pose-guided 2D feature $\mathcal{F}$ to the 3D feature $\mathcal{F}_{3D}$, then extend the 2D features by incorporating 3D joint feature sampling. To integrate and align 2D and 3D features, we utilize a fusion transformer  (Li et al., 2023b) to regress SMPL parameters. Moreover, we employ a pre-trained VQVAE (Van Den Oord et al., 2017), which is trained on a large-scale motion dataset AMASS (Mahmood et al., 2019) with extensive SMPL pose parameters to provide adequate human pose priors and preserve the correspondence of the VQGAN framework, which can obtain discrete representations of human poses. During the regression, the decoder of the VQVAE is utilized to get the pose parameters $\Theta$, while the shape parameters $\beta$ and camera parameters $c$ are directly predicted using linear layers.
@@ 4.4. Objective Function
Character bones often appear longer than their actual length because of imprecision in drawings or the use of artistic interpretation  (Hogarth, 1970; Johnston and Thomas, 1981; Stanchfield, 2007) in human drawings. Due to inconsistent representations of bone length, directly utilizing absolute joint positions becomes impractical. And art literature has consistently highlighted the importance of accurately describing joint angles. Therefore, we expect the 3D bone projections to align with the bones depicted in 2D, ensuring that the reconstructed 3D joint angles have corresponding projections to the depicted 2D joint angles.
@@
PyMAF (Zhang et al., 2021) 312.7 299.4 187.5 301.5 291.2 187.0 143.1 117.4 101.3
EFT (Joo et al., 2021) 144.6 144.4 98.9 168.8 158.9 103.4 158.7 133.1 111.6
HybrIK (Li et al., 2021) 348.7 345.5 199.6 365.2 352.0 208.5 211.0 177.7 144.0
CLIFF (Li et al., 2022) 186.4 181.5 142.2 217.3 201.3 144.5 137.3 113.3 98.0
HMR2.0 (Goel et al., 2023) 118.3 105.0 85.1 181.4 151.4 107.4 128.0 104.6 88.1
MotionBERT (Zhu et al., 2023) 170.6 165.2 120.1 189.1 172.2 127.6 124.4 99.4 83.6
DPMesh (Zhu et al., 2024b) 130.6 121.5 95.3 169.9 152.1 103.2 123.4 98.0 81.1
DPMesh(Retrained) (Zhu et al., 2024b) 127.7 121.4 94.1 166.4 147.1 94.3 122.6 97.3 80.6
Sketch2Pose (Brodt and Bessmeltsev, 2022) 103.8 101.4 78.1 145.5 135.9 86.8 152.1 125.9 100.3
@@
Empirically, artists typically do not rely on exact mathematical measurements for orthographic or perspective projections when creating drawings (Hogarth, 1970; Stanchfield, 2007). Thus, directly reconstructing 3D poses from predicted 2D poses frequently results in highly inaccurate estimations of the angles formed between the bones and the screen. For bone $i$, the angle between the character in the drawing and the screen can be represented as the angle between the 3D vector of the skeleton and the 2D vector of its projection. The foreshortening loss for the skeleton can thus be formulated as:
@@
Self-contacts are prevalent in common human poses, which are revealed by prior works (Hogarth, 1970). We hypothesize that human observers often rely on perceived self-contacts to solve the problem of depth ambiguity and link touching body parts to similar depths. Previous works focus on optimizing regions based on manually annotated self-contact areas. They enforce physical contact between pairs of vertices by mapping each contact region onto the vertices of the roughly aligned SMPL mesh. In contrast, our dataset includes accurate SMPL pose parameters for the human body in sketches, which provides correct relative depth and joint positions of the character skeleton. In our method, we replace the previous self-contact loss with the SMPL pose parameter loss, which is calculated by $L_{1}$ loss between the predicted SMPL pose parameter and the ground-truth SMPL pose parameter, thus supervising the human mesh to recover the correct positions.
@@ 5.1. Implementation Details
• The artist-designed dataset is provided by Sketch2Pose (Brodt and Bessmeltsev, 2022), which contains six sketches with corresponding 3D poses manually modeled by two artists that best align with the artist’s intentions. The merit of this validation set lies in its accurate representation of the ideal 3D pose intended by the artists, whereas its limitation is the scant data volume, comprising merely six very challenging poses.
@@
We train three models for data creation and sketch-to-pose prediction: ControlNet to synthesize sketch data, ViTPose for 2D keypoint detection, and our core sketch-to-pose network. For ControlNet, condition maps and sketches are padded to $512\times 512$. Then, BLIP2 (Li et al., 2023a) is leveraged to generate prompt labels for the sketches. ViTPose is fine-tuned separately on our collected and synthesized sketch data and remains frozen during core-model training, serving solely as a 2D joints extractor. The core model training process consists of two stages. All training and experiments run on 4 Nvidia A6000 GPUs. The supplementary material provides more training details.

### v2

@@ 4.2. Sketch Feature Extractor
It is also known that ControlNet leverages trainable copies of the encoding layers within the denoising U-Net, serving as a robust backbone to learn various conditional controls, significantly enhancing the fine-grained spatial controllability of the Latent Diffusion Model (LDM)  (rombach2022high). In our implementation, we utilize the ControlNet architecture to handle pose-conditioned information from the 2D guidance $\mathcal{G}$ and integrate it into the image features within the denoising U-Net $\boldsymbol{\epsilon_{\theta}}$. The output $\mathcal{F}$ in the decoding layers of $\boldsymbol{\epsilon_{\theta}}$ is expressed as:
@@ 4.3. SMPL Mesh Regressor
In the last stage, an SMPL mesh regressor is proposed to predict 3D poses from the previously extracted features $\mathcal{F}$. Specifically, $p_{\phi_{3}}(\boldsymbol{y}\!\mid\!\mathcal{F})$ refers to the prediction head that generates parameters of the body model from the hierarchical feature maps $\mathcal{F}$. We first lift the pose-guided 2D feature $\mathcal{F}$ to the 3D feature $\mathcal{F}_{3D}$, then extend the 2D features by incorporating 3D joint feature sampling. To integrate and align 2D and 3D features, we utilize a fusion transformer  (li2023jotr) to regress SMPL parameters. Moreover, we employ a pre-trained VQVAE (van2017neural), which is trained on a large-scale motion dataset AMASS (mahmood2019amass) with extensive SMPL pose parameters to provide adequate human pose priors and preserve the correspondence of the VQGAN framework, which can obtain discrete representations of human poses. During the regression, the decoder of the VQVAE is utilized to get the pose parameters $\Theta$, while the shape parameters $\beta$ and camera parameters $c$ are directly predicted using linear layers.
@@ 4.4. Objective Function
Character bones often appear longer than their actual length because of imprecision in drawings or the use of artistic interpretation  (hogarth1970dynamic; johnston1981illusion; stanchfield2007gesture) in human drawings. Due to inconsistent representations of bone length, directly utilizing absolute joint positions becomes impractical. And art literature has consistently highlighted the importance of accurately describing joint angles. Therefore, we expect the 3D bone projections to align with the bones depicted in 2D, ensuring that the reconstructed 3D joint angles have corresponding projections to the depicted 2D joint angles.
@@
PyMAF (zhang2021pymaf) 312.7 299.4 187.5 301.5 291.2 187.0 143.1 117.4 101.3
EFT (joo2020eft) 144.6 144.4 98.9 168.8 158.9 103.4 158.7 133.1 111.6
HybrIK (li2021hybrik) 348.7 345.5 199.6 365.2 352.0 208.5 211.0 177.7 144.0
CLIFF (li2022cliff) 186.4 181.5 142.2 217.3 201.3 144.5 137.3 113.3 98.0
HMR2.0 (goel2023humans) 118.3 105.0 85.1 181.4 151.4 107.4 128.0 104.6 88.1
MotionBERT (zhu2023motionbert) 170.6 165.2 120.1 189.1 172.2 127.6 124.4 99.4 83.6
DPMesh (zhu2024dpmesh) 130.6 121.5 95.3 169.9 152.1 103.2 123.4 98.0 81.1
DPMesh(Retrained) (zhu2024dpmesh) 127.7 121.4 94.1 166.4 147.1 94.3 122.6 97.3 80.6
Sketch2Pose (brodt2022sketch2pose) 103.8 101.4 78.1 145.5 135.9 86.8 152.1 125.9 100.3
@@
Empirically, artists typically do not rely on exact mathematical measurements for orthographic or perspective projections when creating drawings (hogarth1970dynamic; stanchfield2007gesture). Thus, directly reconstructing 3D poses from predicted 2D poses frequently results in highly inaccurate estimations of the angles formed between the bones and the screen. For bone $i$, the angle between the character in the drawing and the screen can be represented as the angle between the 3D vector of the skeleton and the 2D vector of its projection. The foreshortening loss for the skeleton can thus be formulated as:
@@
Self-contacts are prevalent in common human poses, which are revealed by prior works (hogarth1970dynamic). We hypothesize that human observers often rely on perceived self-contacts to solve the problem of depth ambiguity and link touching body parts to similar depths. Previous works focus on optimizing regions based on manually annotated self-contact areas. They enforce physical contact between pairs of vertices by mapping each contact region onto the vertices of the roughly aligned SMPL mesh. In contrast, our dataset includes accurate SMPL pose parameters for the human body in sketches, which provides correct relative depth and joint positions of the character skeleton. In our method, we replace the previous self-contact loss with the SMPL pose parameter loss, which is calculated by $L_{1}$ loss between the predicted SMPL pose parameter and the ground-truth SMPL pose parameter, thus supervising the human mesh to recover the correct positions.
@@ 5.1. Implementation Details
• The artist-designed dataset is provided by Sketch2Pose (brodt2022sketch2pose), which contains six sketches with corresponding 3D poses manually modeled by two artists that best align with the artist’s intentions. The merit of this validation set lies in its accurate representation of the ideal 3D pose intended by the artists, whereas its limitation is the scant data volume, comprising merely six very challenging poses.
@@
We train three models for data creation and sketch-to-pose prediction: ControlNet to synthesize sketch data, ViTPose for 2D keypoint detection, and our core sketch-to-pose network. For ControlNet, condition maps and sketches are padded to $512\times 512$. Then, BLIP2 (li2023blip) is leveraged to generate prompt labels for the sketches. ViTPose is fine-tuned separately on our collected and synthesized sketch data and remains frozen during core-model training, serving solely as a 2D joints extractor. The core model training process consists of two stages. All training and experiments run on 4 Nvidia A6000 GPUs. The supplementary material provides more training details.
@@ Acknowledgements.
Figure 6. Qualitative Comparison of Multiple Sketch Styles. Our proposed model accurately predicts real human body proportions in cartoon images and outperforms other methods in various sketch styles. Its high performance across multiple sketch styles is attributed to our three-stage pose prediction network design and diverse dataset with perturbations. The red dashed box highlights the unreasonable 3D human pose estimation.
Figure 7. Qualitative Comparison of Challenging Poses. Our method successfully recovers plausible poses even when character parts are obscured or lines overlap, as illustrated in specific scenarios. These results demonstrate that our approach better meets artists’ needs and predicts 3D human poses with higher accuracy. The red dashed box highlights the unreasonable 3D human pose estimation. And our results can be seamlessly applied to a custom character using standard tools.
Figure 8. Results on Sketch Video Input. The frame-by-frame prediction results show our method’s generalization capability in extracting human poses from sketch videos.